OpenStudy (princesssleelee):

If a football player passes a football from 4 feet off the ground with an initial velocity of 36 feet per second, how long will it take the football to hit the ground? Use the equation h = −16t2 + 36t + 4. quantity of 9 plus or minus square root of 65 all over 8 9 plus or minus square root of 65 end root over 8 quantity of 9 plus or minus square root of 97 all over 8 9 plus or minus square root of 97 end root over 8

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle h=-16t^2+36t+4 }\) Where; t is time in seconds. h is height in feet. So, you know that \(\color{#000000 }{ \displaystyle h=4 }\) (height of 4\(\tiny~\)ft).

OpenStudy (princesssleelee):

t=2.36, correct?

OpenStudy (solomonzelman):

Can you please show me how you obtained that value/solution for t?

OpenStudy (princesssleelee):

h=−16t2+36t+4 0=−16t2+36t+4 0=−4t2+9t+1 Using the quadratic formula, we get two values for t: -0.11 and 2.36. Having a negative time doesn't make sense - therefore the answer is 2.36 seconds

OpenStudy (solomonzelman):

Yes

OpenStudy (princesssleelee):

im not sure what to do next.

OpenStudy (princesssleelee):

@SolomonZelman

OpenStudy (princesssleelee):

dont i use the quadratic formula and solve?

OpenStudy (princesssleelee):

do*